How many atoms of sulfur are present in 2.40 moles of sulfur trioxide SO3


1
Expert's answer
2021-09-19T00:03:56-0400

1 mole of SO3 = 6.022 ×1023 molecules (Avogadro's constant)


2.40 moles of SO3 = (2.40) × (6.022 × 1023 ) molecules


=14.4528 × 1023 molecules

In 1 molecule of SO3 there is 1 atom


Therefore in 2.40 moles of SO3 there are also


=14.4528 ×1023 atoms